Using laws of exponents, simplify and write the answer in exponential form: (5^2 )^ 3 ÷ 5^ 3

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

By the rule of taking power of as power = (a^m)^n= a^mn
(5^2)^3can be written as = (5)^2 × 3
​ = 5^6
Now , 5^6÷5^3
By the rule of dividing the powers with same base = a^m÷ a^n= a^m - n
Then ,
= 5^6-3
= 5^3

Thanks